OpenStudy (waleed_imtiaz):

A hollow metallic sphere of 8cm diameter is chatged with 4x10^-8 Coulombs. The potential on its surface will be: A) 900 volts B) 9000 volts C) 90 volts D) zero Please, give me the reason......

OpenStudy (anonymous):

D- because there is no electric field inside the hollow sphere. All the points on the surface are at the same potential. There in no difference in the potential, so no potential difference.

OpenStudy (shane_b):

It's asking for the potential on the surface...not inside the sphere. So it's not D.

OpenStudy (shane_b):

The potential outside the sphere is\[\frac{kq}{r}=\frac{(9x10Nm^2/C^2)(4x10^{-8C})}{0.04m}=9000V\]

OpenStudy (vincent-lyon.fr):

Potential on the surface is equal to potential at the centre. Using U = K Q/r gives 9000 V.

OpenStudy (anonymous):

oh right, just read the question again :)

OpenStudy (waleed_imtiaz):

yes, correct.... I am asking about the potential on the surface...

OpenStudy (shane_b):

When I said outside, I meant outer surface

OpenStudy (waleed_imtiaz):

yes.,.. I got it...... thanks guyz...

OpenStudy (waleed_imtiaz):

yes I understand U @Shane_B

OpenStudy (shane_b):

ok :)

OpenStudy (waleed_imtiaz):

I have a few more questions........

OpenStudy (shane_b):

Post the separately and I'm sure someone will answer them...I have grass to cut :/

OpenStudy (waleed_imtiaz):

oh..... Okay.... No problemo xD